# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v4

@@ V-F Urban and forest LiDAR-only metric-semantic SLAM with inter-robot loop closures
The quantitative results on inter-robot localization are shown in Table V. In urban environments, our algorithm utilizes semantic object models of various classes and shapes to estimate the relative 3D position and yaw between each pair of robots. In forest settings, our method leverages models of tree trunks for inter-robot localization. The SlideMatch algorithm achieves an average position error of 0.13 m and a yaw error of 0.68, with standard deviations of 0.07 m and 1.76, respectively. In contrast, the SlideGraph algorithm shows an average position error of 0.24 m and a yaw error of -0.57, with standard deviations of 0.28 m and 1.60, respectively.
@@ V-G Indoor RGBD-only aerial-ground metric-semantic SLAM with inter-robot loop closures
The quantitative results are shown in Table V. For the indoor experiment with multiple Scarab UGVs, our SlideMatch algorithm achieves an average position error of 0.28 m and a yaw error of 0.10°, with standard deviations of 0.13 m and 2.90°, respectively. Our SlideGraph algorithm achieves an average position error of 0.75 m and a yaw error of -2.77°, with standard deviations of 0.35 m and 2.55°, respectively.
Fig. 9 shows the results constructed by the Falcon 250 and one of the UGVs that autonomously explore the environment. The invariance of semantic object models to viewpoint changes enables our algorithm to merge measurements from different viewpoints as observed by heterogeneous aerial and ground robots. As shown in the last row of Table V, the SlideMatch algorithm successfully identifies inter-robot loop closures and estimates the relative transformation with a position error of 0.245 m and a yaw error of 2.35. However, we observed that aerial-ground localization exhibits slightly larger errors along the Z-axis compared to localization errors among identical robot platforms. This discrepancy arises because aerial robots primarily observe the upper parts of objects, while ground robots focus on the lower parts, leading to greater variation in Z-axis estimates. In contrast, the SlideGraph algorithm fails to perform inter-robot loop closures between RGBD-based aerial and ground robots. This limitation is caused by reduced pairwise consistency between object positions in the map, which stems from noisy RGB-D sensing combined with the fact that the aerial and ground robots observe different parts of the objects.
@@ V-H Indoor LiDAR-RGBD metric-semantic SLAM with inter-robot loop closures
Quantitative results are shown in Table V. Our semantics-driven place recognition algorithms demonstrate the ability to establish loop closures between robots equipped with different sensors. SlideMatch estimates the relative transformation with a position error of 0.542 m and a yaw error of -4.00, while SlideGraph results in a position error of 0.578 m and a yaw error of 2.00. The accuracy of inter-robot localization in cross-sensing-modality scenarios is slightly lower compared to single-modality experiments, particularly those involving LiDAR-based robots. This performance decline can be attributed to differences in fields of view, object detection accuracy, and noise levels between LiDAR and RGBD sensors. These factors lead to greater discrepancies in object modeling, which, in turn, reduce the precision of inter-robot localization.
@@ V-I Object mapping quantitative results
We report precision, recall, and F1 scores on object mapping results of both cars and tree trunks across multiple experiments. We first obtain the ground truth by surveying the environment and manually marking down the positions of objects. Here, we focus the object mapping results on cars and tree trunks since ground truth for them can be reliably acquired when compared to other objects such as chairs, tables or monitors. This is because cars are usually parked in clearly separated parking spaces and tree trunks are often spaced apart at regular intervals. A True Positive (TP), or a valid object match, occurs when an object in our metric-semantic map is confirmed to exist at a position that aligns with the ground truth. A False Positive (FP) refers to an object that appears in our metric-semantic map but is not present in the ground truth. Conversely, a False Negative (FN) refers to an object that is present in the ground truth but does not appear in our metric-semantic map. These values are then used to calculate precision, recall, and F1 scores. The results can be found in Table VI. Note that those results are acquired using the smallest backbone (darknet-13 as detailed in [3]) for the semantic segmentation model, which is less accurate but can run inference in real time using the onboard NUC computer. The results can be further improved using larger backbones such as darknet-26 or darknet-53, which are supported in [55]. As shown Table VI, the average precision for cars is 0.967 and for tree trunks is 0.946. The average recall for cars is 0.881 and for tree trunks is 0.909. The F1 score is 0.922 for cars and 0.927 for tree trunks. In summary, through these metrics, we demonstrate that our method is able to reliably detect and map objects that exist in the environment, with very few false positives or false negatives. This result is important for downstream tasks such as inventory estimation.
@@
Table VI: Quantitative results of object detection: Ground Truth (G.T.) vs. Estimated (Est.). Precision is evaluated by determining the proportion of detected objects that accurately match the ground truth, thus measuring the algorithm’s object mapping accuracy. Recall indicates the likelihood of the algorithm successfully detecting an object present in the environment, reflecting the algorithm’s object mapping sensitivity.
@@ V-J Semantic KITTI results
Qualitative results, presented in Fig. 7, demonstrate the ability of the proposed algorithm framework to construct metric-semantic maps and merge measurements from multiple robots. We conducted a quantitative evaluation of our SlideGraph algorithm for inter-robot loop closure using sequences #5 and #7. Sequence #7 was divided into three intervals, i.e., (0, 600), (400, 1000), and (500, 1100), for the first, second, and third robots, respectively. Sequence #5 was similarly divided into three intervals, i.e., (0, 1000), (700, 2000), and (1600, 2760), to create datasets for each of the three robots. Our algorithm successfully identified inter-robot loop closures and accurately estimated the relative transformations between the robots. For sequence #5, the SlideGraph algorithm results in a position error of 0.03 m and 0.05 m, and a rotation error of 0.080 and 0.065 for the transformations between robot 0 and robot 1, and between robot 1 and robot 2, respectively. In sequence #7, the algorithm results in a position error of 0.093 m and 0.095 m, and a rotation error of 0.096 and 0.097, respectively.
@@ V-K2 Experimental results
In this section, we show the computational and communication usage when running our metric-semantic SLAM framework. All results are obtained using a computer equipped with an Intel i7-10750H CPU, without the support of a GPU. This CPU’s performance is comparable to that of the onboard computers in the robots. In Table VII, the average and maximum size of communication packets are summarized when robots operate with intermittent communication. To replicate intermittent communication, each robot publishes ROS messages to each other every 10 seconds. Benefiting from the sparse metric-semantic map representation, our algorithm is shown to be communication efficient when deployed on real robots. The similarity in communication usage across different robots stems from the fact that we replicate all robots as being within communication range by streaming their data on different ROS nodes running on a laptop. For indoor experiments, communication is facilitated via Wi-Fi. Although each robot shares all the information in its database with neighboring robots within the communication range, we can further reduce the communication burden by implementing a bookmarking system that only publishes new information since the last communication between robots.
@@
Outdoor Forest 860 0 326.476 28.702 1091 / 1124 1.702 2.686
1 367.798 28.638 980 / 2091 1.735 2.687
2 746.870 52.046 1491 / 1836 1.742 2.686
Table VII: Runtime and communication results. The number of objects is the count of object landmarks in our metric-semantic map. Factor adding and graph optimization include both exploration (i.e., observing new landmarks) and loop closure scenarios. The runtime on semantics-driven place recognition and loop closure algorithms includes both loop closure detection and relative transformation estimation. Note that the place recognition and loop closure process runs on a separate thread. Therefore, its delay does not impact real-time operation. SlideGraph requires significantly lower computational runtime, particularly in environments with fewer object landmarks. This efficiency is attributed to the use of descriptor-based matching, where computational demands are influenced not by the scale of the environment but only by the number of object landmarks. In contrast, SlideMatch exhaustively searches across the entire search region, and its runtime depends on various factors, such as the search discretization. The communication usage is measured by the size of messages published by each robot. Here, the average usage is calculated by dividing the total size of all messages by the number of messages, while the maximum size corresponds to the size of the last messages since they usually contain the most observations.
Additionally, the runtime for each module, including object detection and modeling, factor adding and graph optimization, and loop closure (if applicable), is reported in Table VII. Except for the loop closure module, which is used to identify inter-robot loop closures and operates on a separate CPU thread, all other modules are triggered each time new observations arrive.
@@
These results illustrate that the average runtime required for factor graph adding and graph optimization increases with the number of landmarks. This relationship is likely influenced by the presence of one or more inter-robot loop closures in these experiments, which makes the computational complexity lie between pure exploration (observing new landmarks) and processing loop closures.
@@
In addition to the importance of sparse semantic landmarks in facilitating long-range autonomous exploration and navigation, our results on communication provide another reason for using semantic landmarks in multi-robot collaboration tasks. The average communication per landmark is 3.96 kilobytes (KB) for indoor environments, 1.84 KB for outdoor environments, and 2.03 KB for forest environments. The average communication bandwidth per meter traveled by the robot is 1.65 KB in indoor settings, 1.85 KB in outdoor settings, and 5.13 KB in forest settings. As the density of landmarks in the environment increases, communication overhead also increases. However, the overall communication bandwidth required remains minimal. This efficiency in communication shows the effectiveness of sparse semantic mapping in minimizing resource usage.
@@ VI Discussion
SlideGraph demonstrates superior performance with minimal parameter tuning in scenarios where the object-level metric-semantic map experiences relatively low noise in object position estimates. Therefore, SlideGraph is generally the preferred option, particularly for LiDAR-based experiments or scenarios involving large search spaces. On the other hand, the SlideMatch algorithm can sometimes achieve successful loop closure in situations where the SlideGraph method fails, particularly in experiments with large errors in object position estimates. This is because the pairwise consistency assumption used in [31] becomes less suitable under such conditions. Such situations are common in indoor RGBD-based object mapping, where cluttered spaces lead to small differences in object-to-object distances and high noise in position estimates due to noisy depth measurements, noisy local odometry, limited field of view, and complex object shapes. Although semantic labels can help disambiguate these measurements to some extent, it remains challenging to use pairwise consistency to distinguish outliers from inliers. In contrast, SlideMatch leverages the entire map and functions more like a voting method, making it more resilient to noise in local pairwise object distances.
VII Conclusion
@@
In this work, we introduce a real-time decentralized metric-semantic SLAM framework, along with its integration into autonomous exploration and navigation systems for aerial and ground robots. This integrated system enables teams of heterogeneous robots to autonomously explore and collaboratively construct maps with both geometric and open-vocabulary semantic information across a variety of indoor and outdoor environments. Each robot independently executes the metric-semantic SLAM framework and the autonomous exploration stack onboard in real time, operating in a decentralized manner. The robots opportunistically leverage communication to exchange sparse, lightweight semantic measurements, which are used for inter-robot localization and map merging. Through a comprehensive set of real-world experiments involving three types of aerial and ground robots, we demonstrate the capabilities, effectiveness, and robustness of our system. Quantitative results on metrics such as inter-robot localization accuracy and object mapping precision from these experiments, as well as additional experiments conducted on the public Semantic KITTI dataset, demonstrate the accuracy and robustness of the proposed framework. Our runtime and communication analysis also shows the importance of using sparse semantic map representations, especially for large-scale metric-semantic mapping tasks in challenging and dense environments. For instance, the entire database containing poses and object observations from a mission covering a 1-km trajectory with 1,000 object landmarks amounts to merely 2-3 megabytes. We envision that the proposed metric-semantic SLAM framework, along with its integration and deployment on autonomous aerial and ground robots, will pave the way for future research in semantically informed robotic perception and decision-making as well as multi-robot collaboration.
VIII Acknowledgment

### v5

@@ V-F Urban and forest LiDAR-only metric-semantic SLAM with inter-robot loop closures
The quantitative results on inter-robot localization are shown in Table IV. In urban environments, our algorithm utilizes semantic object models of various classes and shapes to estimate the relative 3D position and yaw between each pair of robots. In forest settings, our method leverages models of tree trunks for inter-robot localization. The SlideMatch algorithm achieves an average position error of 0.13 m and a yaw error of 0.68, with standard deviations of 0.07 m and 1.76, respectively. In contrast, the SlideGraph algorithm shows an average position error of 0.24 m and a yaw error of -0.57, with standard deviations of 0.28 m and 1.60, respectively.
@@ V-G Indoor RGBD-only aerial-ground metric-semantic SLAM with inter-robot loop closures
The quantitative results are shown in Table IV. For the indoor experiment with multiple Scarab UGVs, our SlideMatch algorithm achieves an average position error of 0.28 m and a yaw error of 0.10°, with standard deviations of 0.13 m and 2.90°, respectively. Our SlideGraph algorithm achieves an average position error of 0.75 m and a yaw error of -2.77°, with standard deviations of 0.35 m and 2.55°, respectively.
Fig. 9 shows the results constructed by the Falcon 250 and one of the UGVs that autonomously explore the environment. The invariance of semantic object models to viewpoint changes enables our algorithm to merge measurements from different viewpoints as observed by heterogeneous aerial and ground robots. As shown in the last row of Table IV, the SlideMatch algorithm successfully identifies inter-robot loop closures and estimates the relative transformation with a position error of 0.245 m and a yaw error of 2.35. However, we observed that aerial-ground localization exhibits slightly larger errors along the Z-axis compared to localization errors among identical robot platforms. This discrepancy arises because aerial robots primarily observe the upper parts of objects, while ground robots focus on the lower parts, leading to greater variation in Z-axis estimates. In contrast, the SlideGraph algorithm fails to perform inter-robot loop closures between RGBD-based aerial and ground robots. This limitation is caused by reduced pairwise consistency between object positions in the map, which stems from noisy RGB-D sensing combined with the fact that the aerial and ground robots observe different parts of the objects.
@@ V-H Indoor LiDAR-RGBD metric-semantic SLAM with inter-robot loop closures
Quantitative results are shown in Table IV. Our semantics-driven place recognition algorithms demonstrate the ability to establish loop closures between robots equipped with different sensors. SlideMatch estimates the relative transformation with a position error of 0.542 m and a yaw error of -4.00, while SlideGraph results in a position error of 0.578 m and a yaw error of 2.00. The accuracy of inter-robot localization in cross-sensing-modality scenarios is slightly lower compared to single-modality experiments, particularly those involving LiDAR-based robots. This performance decline can be attributed to differences in fields of view, object detection accuracy, and noise levels between LiDAR and RGBD sensors. These factors lead to greater discrepancies in object modeling, which, in turn, reduce the precision of inter-robot localization.
@@ V-I Object mapping quantitative results
We report precision, recall, and F1 scores on object mapping results of both cars and tree trunks across multiple experiments. We first obtain the ground truth by surveying the environment and manually marking down the positions of objects. Here, we focus the object mapping results on cars and tree trunks since ground truth for them can be reliably acquired when compared to other objects such as chairs, tables or monitors. This is because cars are usually parked in clearly separated parking spaces and tree trunks are often spaced apart at regular intervals. A True Positive (TP), or a valid object match, occurs when an object in our metric-semantic map is confirmed to exist at a position that aligns with the ground truth. A False Positive (FP) refers to an object that appears in our metric-semantic map but is not present in the ground truth. Conversely, a False Negative (FN) refers to an object that is present in the ground truth but does not appear in our metric-semantic map. These values are then used to calculate precision, recall, and F1 scores. The results can be found in Table V. Note that those results are acquired using the smallest backbone (darknet-13 as detailed in [3]) for the semantic segmentation model, which is less accurate but can run inference in real time using the onboard NUC computer. The results can be further improved using larger backbones such as darknet-26 or darknet-53, which are supported in [56]. As shown Table V, the average precision for cars is 0.967 and for tree trunks is 0.946. The average recall for cars is 0.881 and for tree trunks is 0.909. The F1 score is 0.922 for cars and 0.927 for tree trunks. In summary, through these metrics, we demonstrate that our method is able to reliably detect and map objects that exist in the environment, with very few false positives or false negatives. This result is important for downstream tasks such as inventory estimation.
@@
Table V: Quantitative results of object detection: Ground Truth (G.T.) vs. Estimated (Est.). Precision is evaluated by determining the proportion of detected objects that accurately match the ground truth, thus measuring the algorithm’s object mapping accuracy. Recall indicates the likelihood of the algorithm successfully detecting an object present in the environment, reflecting the algorithm’s object mapping sensitivity.
@@ V-J Semantic KITTI results
The qualitative results, as shown in Fig. 7, demonstrate the ability of the proposed framework to construct metric-semantic maps and merge measurements from multiple robots. We conducted a quantitative evaluation of our SlideGraph algorithm for inter-robot loop closure using sequences #$\#5$ and #$\#7$. Sequence #$\#7$ was divided into three intervals, i.e., (0, 600), (400, 1000), and (500, 1100), for the first, second, and third robots, respectively. Sequence #$\#5$ was similarly divided into three intervals, i.e., (0, 1000), (700, 2000), and (1600, 2760), to create datasets for each of the three robots. Our algorithm successfully identified inter-robot loop closures and accurately estimated the relative transformations between the robots. For sequence #$\#5$, the SlideGraph algorithm results in a position error of 0.03 m and 0.05 m, and a rotation error of 0.080 and 0.065 for the transformations between robot 0 and robot 1, and between robot 1 and robot 2, respectively. In sequence #$\#7$, the algorithm results in a position error of 0.093 m and 0.095 m, and a rotation error of 0.096 and 0.097.
@@ V-K2 Experimental results
In this section, we show the computational and communication usage when running our metric-semantic SLAM framework. All results are obtained using a computer equipped with an Intel i7-10750H CPU, without the support of a GPU. This CPU’s performance is comparable to that of the onboard computers in the robots. In Table VI, the average and maximum size of communication packets are summarized when robots operate with intermittent communication. To replicate intermittent communication, each robot publishes ROS messages to each other every 10 seconds. Benefiting from the sparse metric-semantic map representation, our algorithm is shown to be communication efficient when deployed on real robots. The similarity in communication usage across different robots stems from the fact that we replicate all robots as being within communication range by streaming their data on different ROS nodes running on a laptop. For indoor experiments, communication is facilitated via Wi-Fi. Although each robot shares all the information in its database with neighboring robots within the communication range, we can further reduce the communication burden by implementing a bookmarking system that only publishes new information since the last communication between robots.
@@
Outdoor Forest 860 0 326.476 28.702 209 / 1124 1.702 2.686
1 367.798 28.638 406 / 2091 1.735 2.687
2 746.870 52.046 375 / 1836 1.742 2.686
Table VI: Runtime and communication results. The number of objects is the count of object landmarks in our metric-semantic map. The runtime for factor addition and graph optimization accounts for handling both the observation of new landmarks and loop closure scenarios. The runtime on semantics-driven place recognition and loop closure algorithms includes both loop closure detection and relative transformation estimation. Note that the place recognition and loop closure process runs on a separate thread. Therefore, its delay does not impact real-time operation. SlideGraph requires significantly lower computational runtime, particularly in environments with fewer object landmarks. This efficiency is attributed to the use of descriptor-based matching, where computational demands are influenced not by the scale of the environment but only by the number of object landmarks. In contrast, SlideMatch exhaustively searches across the entire search region, and its runtime depends on various factors, such as the search discretization. The communication usage is measured by the size of messages published by each robot. Here, the average usage is calculated by dividing the total size of all messages by the number of messages, while the maximum size corresponds to the size of the last messages since they usually contain the most observations.
Additionally, the runtime for each module, including object detection and modeling, factor adding and graph optimization, and loop closure (if applicable), is reported in Table VI. Except for the loop closure module, which is used to identify inter-robot loop closures and operates on a separate CPU thread, all other modules are triggered each time new observations arrive.
@@
These results illustrate that the average runtime required for factor graph adding and graph optimization increases with the number of landmarks. This relationship is influenced by the presence of inter-robot loop closures in these experiments, which makes the computational complexity lie between observing new landmarks and processing loop closures.
@@
In addition to the importance of sparse semantic landmarks in facilitating long-range autonomous exploration and navigation, our results on communication provide another reason for using semantic landmarks in multi-robot collaboration tasks. The average communication per landmark is 3.96 kilobytes (KB) for indoor environments, 1.84 KB for outdoor environments, and 2.03 KB for forest environments. The average communication bandwidth per meter traveled by the robot is 1.65 KB in indoor settings, 1.85 KB in outdoor settings, and 5.13 KB in forest settings. As the density of landmarks in the environment increases, communication overhead also increases. However, the overall communication bandwidth required remains minimal. This modest communication demand is particularly important when robots operate under intermittent and bandwidth-limited communication conditions. In addition, this efficiency in communication also shows the effectiveness of sparse semantic mapping in minimizing resource usage.
@@ V-L Analysis of SlideMatch and SlideGraph algorithms
Throughout these experiments, we have observed unique advantages, limitations, and deployment considerations for both the SlideGraph and SlideMatch algorithms. SlideGraph demonstrates superior performance with minimal parameter tuning in scenarios where the object-level metric-semantic map experiences relatively low noise in object position estimates. Therefore, SlideGraph is generally the preferred option, particularly for LiDAR-based experiments or scenarios involving large search spaces. On the other hand, the SlideMatch algorithm can sometimes achieve successful loop closure in situations where the SlideGraph method fails, particularly in experiments with large errors in object position estimates. This is because the pairwise consistency assumption used in [32] becomes less suitable under such conditions. Such situations are common in indoor RGBD-based object mapping, where cluttered spaces lead to small differences in object-to-object distances and high noise in position estimates due to noisy depth measurements, noisy local odometry, limited field of view, and complex object shapes. Although semantic labels can help disambiguate these measurements to some extent, it remains challenging to use pairwise consistency to distinguish outliers from inliers. In contrast, SlideMatch leverages the entire map and operates similarly to a voting method, making it more tolerant to noise in local pairwise object distances. As a result, SlideMatch performs better in such worst-case scenarios, albeit at the cost of significant computational overhead.
VI Conclusion
@@
In this work, we introduce a real-time decentralized metric-semantic SLAM framework, along with its integration into autonomous exploration and navigation systems for aerial and ground robots. This integrated system enables teams of heterogeneous robots to autonomously explore and collaboratively construct maps with both geometric and open-vocabulary semantic information across a variety of indoor and outdoor environments. Each robot independently executes the metric-semantic SLAM framework and the autonomous exploration stack onboard in real time, operating in a decentralized manner. The robots opportunistically leverage communication to exchange sparse, lightweight semantic measurements, which are used for inter-robot localization and map merging. Through a comprehensive set of real-world experiments involving three types of aerial and ground robots, we demonstrate the capabilities, effectiveness, and robustness of our system. Quantitative metrics, such as inter-robot localization errors and object mapping accuracy, along with additional results from experiments on the Semantic KITTI dataset, demonstrate the performance of the proposed framework. Our runtime and communication analysis also shows the importance of using sparse semantic map representations, especially for large-scale metric-semantic mapping tasks in challenging and dense environments. For instance, the entire database containing poses and object observations from a mission covering a 1-km trajectory with 1,000 object landmarks amounts to merely 2-3 megabytes. There are several directions for future work. The first is leveraging the heterogeneity in the robot team for active collaborative localization and mapping, building upon our previous works that actively position LiDAR-equipped UGVs as mobile beacons to improve the state estimation [66] or facilitate inter-robot map merging [67] of vision-driven UAVs. The second direction is combining our framework with advanced multi-robot communication frameworks [68] or resource-efficient multi-robot coordination methods [69]. The third direction involves utilizing our sparse, semantically meaningful, open-vocabulary map representation to guide real-time multi-robot decision-making in open-world settings, enabling tasks such as collaborative object search and manipulation in large-scale environments without relying on infrastructure.
VII Acknowledgment